4)The common stock of Corning Corporation sells for $47.60 a share. The stock is expected to pay $2.10 per share next year when the annual dividend is distributed. Corning has established a pattern of increasing its dividends by 6% annually and expects to continue doing so. What is the market rate of return on this stock?

11.20%

10.83%

10.41%

9.87%

9.45%

5)Dover Corporation announced today that its next annual dividend has been set at $2.20 a share. The company also announced that all future dividends will increase by 7% annually. What is the maximum amount you should pay to purchase a share of Dover's stock if your goal is to earn a 14.2% rate of return?

$30.56

$32.34

$35.59

$27.41

$25.65
Answer #1

3). Required Return = [Expected Dividend / Current Price] + Growth Rate

= [$2.10 / $47.60] + 0.06 = 0.0441 + 0.06 = 0.1041, or 10.41%

Hence, 3rd option is correct.

4). Share Price = Next Dividend / [r - g]

= $2.20 / [0.142 - 0.07] = $2.20 / 0.072 = $30.56

Hence, 1st option is correct.